# Pulsar timing array sensitivity to anisotropies in the gravitational wave background

### v1

@@ Appendix A Approximations in the derivation of the covariance matrix in the frequency domain
In the left panel of Fig. 7 we show the diagonal part of the covariance matrix for Fourier coefficients in Eq. (42) (blue full) compared to the approximation in Eq. (43) (green dashed). Here, we divide by the HD correlation and assume a signal consistent with current PTA observations $\Omega_{\mathrm{GW}}h^{2}=6.3\times 10^{-8}(\mathrm{yr}\times f)^{2}$ Antoniadis:2023ott . We have moreover generalized the expressions in Eq. (42) and in Eq. (21) by replacing $S_{h}\to\mathcal{T}S_{h}$ to include the transmission function $\mathcal{T}(f)=[1+1/(fT_{\text{obs}})]^{6}$ Babak:2024yhu ; Hazboun:2019vhv to model the loss in sensitivity at low frequency and make the covariance matrix finite for the assumed spectrum. The transmission function causes the slight dip at $T_{\text{obs}}f_{k}=1$. The approximation differs from the full result by $\sim 1-16\%$ for non-zero frequencies. In the right panel of Fig. 7 we show the full covariance matrix for the first 10 non-zero frequencies. We clearly see that the diagonal dominates, but the off-diagonal components can reach $1-10\%$. This indicates that the approximation taking the covariance matrix to be diagonal in frequency with entries as given in Eq. (21) is valid at the few-percent level, but deserves further investigation in the future with increased PTA sensitivity. Note however that for less smooth GWB spectra than the power law we assume here, the difference between the approximation and the full result can generally be much larger.
@@ Appendix B Generalised overlap reduction functions
In this appendix, we reproduce analytical expressions for the generalized ORFs in the so-called “computational frame”, where the pulsar direction $\hat{p}_{I}$ is aligned in the $\hat{z}$ axis $\hat{p}_{I}=(0,0,1)$ while the second pulsar is placed in the $(\hat{x},\hat{z})$ plane, i.e., $\hat{p}_{J}=(\sin\zeta,0,\cos\zeta)$ Mingarelli:2013dsa . Higher multipoles can be found at Gair:2014rwa ; Kato:2015bye ; AnilKumar:2023yfw .
@@
These expressions are visualized in Fig. 8. We note in particular the sizeable hierarchy between the values at $\zeta=0$ and $\zeta\neq 0$ for the monopole and dipole, which leads to a particularly good sensitivity of PTAs to these low multipoles.
Figure 8: Real-valued ORF in the computational frame as a function of the pular-pular angular separation $\zeta$. We include pulsar auto-correlations. The dots report the ORFs between a pulsar located along $\hat{z}$ and a random set of 100 pulsars in the $(\hat{x},\hat{z})$ plane, computed numerically by projecting Eq. (13) on the spherical harmonics. The dashed lines indicate the analytical results in Eqs. (48), (49), and (50).

### v2

@@ Appendix A Approximations in the derivation of the covariance matrix in the frequency domain
In the left panel of Fig. 7 we show the diagonal part of the covariance matrix for Fourier coefficients in Eq. (47) (blue full) compared to the approximation in Eq. (50) (green dashed). Here, we divide by the HD correlation and assume a signal consistent with current PTA observations $\Omega_{\mathrm{GW}}h^{2}=6.3\times 10^{-8}(\mathrm{yr}\times f)^{2}$ Antoniadis:2023ott . We have moreover generalized the expressions in Eq. (47) and in Eq. (21) by replacing $S_{h}\to\mathcal{T}S_{h}$ to include the transmission function $\mathcal{T}(f)=[1+1/(fT_{\text{obs}})]^{6}$ Babak:2024yhu ; Hazboun:2019vhv to model the loss in sensitivity at low frequency and make the covariance matrix finite for the assumed spectrum. The transmission function causes the slight dip at $T_{\text{obs}}f_{k}=1$. The approximation differs from the full result by $\sim 1-16\%$ for non-zero frequencies. In the right panel of Fig. 7 we show the full covariance matrix for the first 10 non-zero frequencies. We clearly see that the diagonal dominates, but the off-diagonal components can reach $1-10\%$. This indicates that the approximation taking the covariance matrix to be diagonal in frequency with entries as given in Eq. (21) is valid at the few-percent level, but deserves further investigation in the future with increased PTA sensitivity. Note however that for less smooth GWB spectra than the power law we assume here, the difference between the approximation and the full result can generally be much larger. See also Allen:2024uqs ; Allen:2024bnk for a discussion on the correlations between frequency bins.
As a final comment, we stress that in the strong signal limit discussed in section IV, even if present, correlations between frequencies do not impact the determination of the $c_{\ell m}$ coefficients. This follows from the signal and FIM structure in the frequency and $\ell m$ spaces. Since the $c_{\ell m}$ coefficients are frequency-independent, the $M$ matrix appearing in Eq. (30) multiplies the identity in the frequency space and the $N_{f}$ scaling of the FIM is unaffected.
@@ Appendix B Generalised overlap reduction functions
In this appendix, we report the analytical expressions for the generalized ORFs in the so-called “computational frame” Mingarelli:2013dsa ; Gair:2014rwa ; Kato:2015bye ; AnilKumar:2023yfw ; Allen:2024bnk . This corresponds to aligning the $\hat{z}$ axis with the first pulsar direction $\hat{p}_{I}=(0,0,1)$, while the second pulsar is placed in the $(\hat{x},\hat{z})$ plane, i.e., $\hat{p}_{J}=(\sin\zeta,0,\cos\zeta)$. Higher multipoles can be found at Gair:2014rwa ; Kato:2015bye ; AnilKumar:2023yfw .
@@
These expressions are visualized in Fig. 8 (and similar figures are shown in Mingarelli:2013dsa ). We note in particular the sizeable hierarchy between the values at $\zeta=0$ and $\zeta\neq 0$ for the monopole and dipole, which leads to a particularly good sensitivity of PTAs to these low multipoles.
Figure 8: Real-valued ORF in the computational frame as a function of the pulsar-pulsar angular separation $\zeta$. We include pulsar auto-correlations. The dots report the ORFs between a pulsar located along $\hat{z}$ and a random set of 100 pulsars in the $(\hat{x},\hat{z})$ plane, computed numerically by projecting Eq. (13) on the spherical harmonics. The dashed lines indicate the analytical results in Eqs. (55), (56), and (57).